# Meta Co-Training: Two Views are Better than One

### v2

@@ Abstract
In many practical computer vision scenarios unlabeled data is plentiful, but labels are scarce and difficult to obtain. As a result, semi-supervised learning which leverages unlabeled data to boost the performance of supervised classifiers have received significant attention in recent literature. One major class of semi-supervised algorithms is co-training. In co-training two different models leverage different independent and sufficient “views” of the data to jointly make better predictions. During co-training each model creates pseudo labels on unlabeled points which are used to improve the other model. We show that in the common case when independent views are not available we can construct such views inexpensively using pre-trained models. Co-training on the constructed views yields a performance improvement over any of the individual views we construct and performance comparable with recent approaches in semi-supervised learning, but has some undesirable properties. To alleviate the issues present with co-training we present Meta Co-Training which is an extension of the successful Meta Pseudo Labels approach to two views. Our method achieves new state-of-the-art performance on ImageNet-10% with very few training resources, as well as outperforming prior semi-supervised work on several other fine-grained image classification datasets.
@@ 1 Introduction
In many practical machine learning scenarios we have access to a large amount of unlabeled data, and relatively few labeled data points for the task on which we would like to train. For generic computer vision (CV) tasks there are large well-known and open source datasets with hundreds of millions or even billions of unlabeled images [62, 63]. In contrast, labeled data is usually scarce and expensive to obtain requiring many human-hours to generate. Hence, in this context, semi-supervised learning (SSL) methods are becoming increasingly popular as they rely on training useful learning models by using small amounts of labeled data and large amounts of unlabeled data.
@@
Not unrelated to, and not to be confused with the idea of semi-supervised learning, is self-supervised learning. [^1] There is also an SSL technique called self-training. Recognizing the unfortunate similarity of these three terms, SSL will always mean semi-supervised learning in this text and we will not abbreviate the other terms. In self-supervised learning a model is trained with an objective that does not require a label that is not evident from the data itself. Self-supervised learning was popularized by the BERT [32] model for generating word embeddings for natural language processing (NLP) tasks. BERT generates its embeddings by solving the pretext task of masked word prediction. Pretext tasks present unsupervised objectives that are not directly related to any particular supervised objective we might desire to solve, but rather are solved in the hope of learning a suitable representation to more easily solve a downstream task. Due to the foundational nature of these models to the success of many models in NLP, or perhaps the fact that they serve as a foundation for other models to be built upon, these pretext learners are often referred to as foundation models. Inspired by foundation models from NLP, several pretext tasks have been proposed for CV with associated foundation models that have been trained to solve them [8, 38, 20, 42, 23, 49, 19, 58, 24, 43]. Unlike NLP, the learned representations for images are often much smaller than the images themselves which yields the additional benefit of reduced computational effort when using the learned representations.
Typically, SSL requires the learner to make predictions on a large amount of unlabeled data as well as the labeled data during training. Furthermore, SSL often requires that the learner be re-trained [12, 76, 50, 72, 48, 55, 39, 70, 74, 40, 17] partially, or from scratch. Given the computational benefits of learning from a frozen representation, combining SSL and self-supervised learning is a natural idea. Additionally, the representation learner allows us to compress our data to a smaller size. This compression is beneficial not only because of the reduction in operations required that is associated with a smaller model, but also because it allows us to occupy less space on disk, in RAM, and in high bandwidth memory on hardware accelerators.
These benefits are particularly pronounced for co-training [12] in which two different “views” of the data must be obtained or constructed and then two different classifiers must be trained and re-trained iteratively. Seldom does a standard problem in machine learning present two independent and sufficient views of the problem to be leveraged for co-training. As a result, if one hopes to apply co-training one usually needs to construct two views from the view that they have. To take advantage of the computational advantages of representation learning, and the large body of self-supervised literature, we propose using the embeddings produced by foundation models trained on different pretext tasks as the views for co-training. To our knowledge we are the first to propose using frozen learned feature representations as views for co-training classification, though we think the idea is very natural. We show that views constructed in this way appear to satisfy the necessary conditions for the effectiveness of co-training.
@@
Unfortunately, while we can boost performance by training different models on two views, co-training performs sub-optimally during its pseudo-labeling step. To more effectively leverage two views we propose a novel method for co-training based on Meta Pseudo Labels [59] called Meta Co-Training in which each model is trained to provide better pseudo labels to the other given only its view of the labeled data and the performance of the other model on the labeled set. We show that our approach provides an improvement over both co-training and the current state-of-the-art (SotA) on the ImageNet-10% dataset, as well as it establishes new SotA few-shot performance on several other fine-grained image classification tasks.
Summary of Contributions.   Our contributions are therefore summarized as follows.
@@
    Use of embeddings obtained from different pre-trained models as different views for co-training.
@@
    We boost the performance of standard co-training by proposing a method that relies on Meta Pseudo Labels [59], called Meta Co-Training.
@@
    We establish new state-of-the-art top-1 accuracy on ImageNet-10% [31] and new state-of-the-art top-1 few-shot accuracy on several other CV classification tasks.
  4. 4.
    We make our implementation publicly available:
    https://github.com/JayRothenberger/Meta-Co-Training
Outline of the Paper.   In Section 2 we discuss related work. In Section 3 we establish notation and provide more information on co-training and meta pseudo labels. In Section 4 we describe our proposed methods of view construction and of meta co-training. In Section 5 we present and discuss our experimental findings on ImageNet [31], Flowers102 [57], Food101 [14], FGVCAircraft [54], iNaturalist [44], and iNaturalist 2021 [45]. In Section 6 we conclude with a summary and ideas for future work.
2 Related Work
While SSL [21, 78, 67, 73] has become very relevant in recent years with the influx of the data collected by various sensors, the paradigm and its usefulness have been studied at least since the 1960’s; e.g., [41, 29, 64, 36]. Among the main approaches for SSL that are not directly related to our work, we find techniques that try to identify the maximum likelihood parameters for mixture models [30, 7], constrained clustering [9, 28, 69, 68, 18, 3], graph-based methods [11, 10, 53, 6], as well as purely theoretical probably approximately correct (PAC)-style results [4, 37, 25]. Below we discuss SSL approaches closer to our work.
Self-Training and Student-Teacher Frameworks.   One of the earliest ideas for SSL is that of self-training [64, 36]. This idea is still very popular with different student-teacher frameworks [79, 76, 59, 50, 72, 66, 55, 39, 17]. This technique is also frequently referred to as pseudo-labeling [48] as the student tries to improve its performance by making predictions on the unlabeled data, then integrating the most confident predictions to the labeled set, and subsequently retraining using the broader labeled set. In addition, one can combine ideas; e.g., self-training with clustering [46], or create ensembles via self-training [47], potentially using graph-based information [53]. An issue with these approaches is that of confirmation bias [2], where the student cannot improve a lot due to incorrect labeling that occurs either because of self-training, or due to a fallible teacher.
Co-Training.   Another paradigmatic family of SSL algorithms are co-training [12] algorithms, where instances have two different views for the same underlying phenomenon. Co-training has been very successful with wide applicability [34] and can outperform learning algorithms that use only single-view data [56]. Furthermore, the generalization ability of co-training is well-motivated theoretically [12, 27, 5, 26]. There are several natural extensions of the base idea, such as methods for constructing different views [71, 15, 22, 65, 16], including connections to active learning [33], methods that are specific to deep learning [60], and other natural extensions that exploit more than two views on data; e.g., [77].
@@ 3 Background
We briefly describe two important components of semi-supervised learning that are used jointly in order to devise our proposed new technique. However, before we do so, we also define basic notions so that we can eliminate ambiguity.
3.1 Notation
Supervised learning is based on the idea of learning a function, called a model, $f$ belonging to some model space $\mathcal{F}$, using training examples $(x,y)$ that exemplify some underlying phenomenon. Training examples are labeled instances; i.e., $x\in\mathcal{X}$ is an instance drawn from some instance space $\mathcal{X}$ and $y$ is a label from some label set $\mathcal{Y}$. This relationship between $\mathcal{X}$ and $\mathcal{Y}$ can be modeled with a probability distribution $D$ that governs $\mathcal{X}\times\mathcal{Y}$. We write $(x,y)\sim D$ to indicate that a training example $(x,y)$ is drawn from $\mathcal{X}\times\mathcal{Y}$ according to $D$. We use $\mathbf{1}\left\{\mathcal{A}\right\}$ as an indicator function of the event $\mathcal{A}$; that is, $\mathbf{1}\left\{\mathcal{A}\right\}$ is equal to 1 when $\mathcal{A}$ holds, otherwise 0. For vectors $\alpha,\beta\in\mathbb{R}^{d}$ we indicate their Hadamard (element-wise) product using $\otimes$; that is, $\gamma=\alpha\otimes\beta$, where $\gamma_{j}=\alpha_{j}\cdot\beta_{j}$ for $j\in\{1,\ldots,d\}$.
As is standard in classification problems, we let model $f$ be a function of the form $f\colon\mathcal{X}\rightarrow\Delta^{|\mathcal{Y}|}$ where $\Delta^{|\mathcal{Y}|}$ is the $|\mathcal{Y}|$-dimensional unit simplex. Equivalently $f$ can be understood as a probability distribution $\mathbf{Pr}(y\ |\ x)$. We use $f(x)|_{j}$ to denote the $j$-th value of $f$ in the output in such a case. It is common for $\mathcal{F}$ to be a parameterized family of functions (e.g. artificial neural networks). We use $\theta$ as a vector that holds such parameters, and we denote the function parameterized by $\theta$ as $f_{\theta}$. An individual prediction $f_{\theta}(x)$ for an instance $x$ is evaluated via the use of a loss function $\ell$ that determines the quality of the prediction $f_{\theta}(x)$ with respect to the actual value $y$ that should have been predicted. For classification problems a natural way of defining loss functions is of the form $\ell\colon\mathcal{Y}\times\Delta^{|\mathcal{Y}|}\rightarrow\mathbb{R}_{+}$. For the entirety of this text $\ell$ will refer to the cross-entropy loss which is defined as $\ell\left(y,f_{\theta}(x)\right)=-\sum_{\xi\in\mathcal{Y}}\mathbf{1}\left\{\xi=y\right\}\log\left(f_{\theta}(x)|_{\xi}\right)$.
We measure the fitness of a model $f_{\theta}$ by its (true) risk $R_{D}\left(f_{\theta}\right)=\mathbf{E}_{(x,y)\sim D}\left[\ell\left(y,f_{\theta}(x)\right)\right]$. While we want to create a model $f_{\theta}$ that has low true risk, in principle we apply empirical risk minimization (ERM) which yields a model $f_{\theta}$ that minimizes the risk over a set of training examples. That is, we use empirical risk as a proxy for the true risk. Given a set $S=\left\{(x_{i},y_{i})\right\}_{i=1}^{\left\lvert S\right\rvert}$, we denote the empirical risk of $f_{\theta}$ on $S$ as $\widehat{R}_{S}\left(f_{\theta}\right)=\frac{1}{\left\lvert S\right\rvert}\sum_{i=1}^{\left\lvert S\right\rvert}\ell\left(y,f_{\theta}(x)\right)$. To ease notation, when the inputs to $\ell$ are a batch of examples $(X,Y)$, we denote $\widehat{R}_{(X,Y)}\left(f_{\theta}\right)$ as simply $\ell(Y,f_{\theta}(X))$. Similarly, when we write the arg max of a batch of vectors we refer to the batch of the arg max of the individual values. In general, ERM can be computationally hard (e.g., [13, 1]); hence, we often rely on gradient-based optimization techniques, such as stochastic gradient descent (SGD), that can take us to a local optimum efficiently. We use $\eta$ as the learning rate and $T$ as the maximum number of steps (or updates) in such a gradient-based optimization procedure.
@@ 3.1 Notation
Semi-Supervised Learning.   We are interested in situations where we have a large pool of unlabeled instances $U$ as well as a small portion $L$ of them that is labeled. In other words, we assume that the learning algorithm has access to a dataset $S=L\cup U$, such that $L=\left\{(x_{i},y_{i})\right\}_{i=1}^{m}$ and $U=\left\{x_{j}\right\}_{j=1}^{u}$. In this setting, semi-supervised (SSL) methods attempt to first learn an initial model $f_{init}$ using the labeled data $L$ via a supervised learning algorithm, and in sequence, an attempt is being made in order to harness the information that is hidden in the unlabeled set $U$, so that a better model $f$ can be learnt.
3.2 Meta Pseudo Labels
The hope is that updated student parameters $\theta_{S}^{\text{PL}}$ will behave well on the labeled data $L=(X_{L},Y_{L})=\left\{(x_{i},y_{i})\right\}_{i=1}^{m}$; i.e., the loss $\mathcal{L}_{L}\left(\theta_{S}^{\text{PL}}\left(\theta_{T}\right)\right)=\ell\left(Y_{L},f_{\theta_{S}^{\text{PL}}}(X_{L})\right)$ is small, where we defined $\mathcal{L}\left(\theta_{S}^{\text{PL}}\left(\theta_{T}\right)\right)$ so that the dependence between $\theta_{S}$ and $\theta_{T}$ is explicit.
However, because the dependency between $\theta_{T}$ and $\theta_{S}$ is complicated, a practical approximation [51, 35] is obtained via $\theta_{S}^{\text{PL}}\approx\theta_{S}-\eta_{S}\cdot\nabla_{\theta_{S}}\mathcal{L}_{u}\left(\theta_{T},\theta_{S}\right)$ which leads to the practical teacher objective:
For more information the interested reader can refer to [59].
3.3 Co-Training
Co-training relies on the existence of two different views of the instance space $\mathcal{X}$; that is, $\mathcal{X}=\mathcal{X}^{(1)}\times\mathcal{X}^{(2)}$. Hence, an instance $x\in\mathcal{X}$ has the form $x=(x^{(1)},x^{(2)})\in\mathcal{X}^{(1)}\times\mathcal{X}^{(2)}$. We say that the views $x^{(1)}$ and $x^{(2)}$ are complementary. Co-training requires two assumptions in order to work. First, it is assumed that, provided enough many examples, each separate view is enough for learning a meaningful model. The second is the conditional independence assumption [12], which states that $x^{(1)}$ and $x^{(2)}$ are conditionally independent given the label of the instance $x=(x^{(1)},x^{(2)})$. Thus, with co-training, two models $f^{(1)}$ and $f^{(2)}$ are learnt, which help each other using different information that they capture from the different views of each instance. More information is presented below.
Given a dataset $S=L\cup U$ (Section 3.1 has notation), we can have access to two different views of the dataset due to the natural partitioning of the instance space $\mathcal{X}$; that is, $S^{(1)}=L^{(1)}\cup U^{(1)}$ and $S^{(2)}=L^{(2)}\cup U^{(2)}$. Initially two models $f^{(1)}$ and $f^{(2)}$ are learnt using a supervised learning algorithm based on the labeled sets $L^{(1)}$ and $L^{(2)}$ respectively. Learning proceeds in iterations, where in each iteration a number of unlabeled instances are integrated with the labeled set by assigning pseudo-labels to them that are predicted by $f^{(1)}$ and $f^{(2)}$, so that improved models $f^{(1)}$ and $f^{(2)}$ can be learnt by retraining on the augmented labeled sets. In particular, in each iteration, each model $f^{(v)}$ predicts a label for each view $x^{(v)}\in U^{(v)}$. The top-$k$ most confident predictions by model $f^{(v)}$ are used to provide a pseudo label for the respective complementary instances and then these pseudo-labeled examples are integrated in to the labeled set for the complementary view. In the next iteration a new model is trained based on the augmented labeled set. Both views of the used instances are dropped from $U$ even if only one of the views may be used for augmenting the appropriate labeled set. This process repeats until the unlabeled sets $U^{(v)}$ are exhausted, yielding models $f^{(1)}$ and $f^{(2)}$ corresponding to the two views. Co-Training is evaluated using the joint prediction of these models which is the re-normalized element-wise product of the two predictions:
Co-training was introduced by Blum and Mitchell [12] and their paradigmatic example was classifying web pages using two different views: (i) the bags of words based on the content that they had and (ii) the bag of words formed by the words that were used for hyperlinks pointing to these webpages. In our context the two different views are obtained via different embeddings that we obtain when we use images as inputs to different foundation models and observe the resulting embeddings.
4 Proposed Methods
4.1 View Construction
Past methods of view construction include manual feature subsetting [33], automatic feature subsetting [22, 65], random feature subsetting [15, 16], random subspace selection [16], and adversarial examples [60]. We propose a new method of view construction which is motivated by modern approaches to semi-supervised and representation learning in computer vision and natural language processing: pretext learning. Given two pretext tasks we can train a model to accomplish each task, then use their learned embedding spaces as separate views for classification. If we have chosen our pretext tasks well, then the representations will be different, compressed, and capture sufficient information to perform the desired downstream task.
While our approach is more widely applicable, it is particularly well-suited to CV. There are a plethora [38, 23, 24, 19, 49, 20, 61, 58, 42] of competitive learned representations for images. Recently it has been popular to treat these representations as task-agnostic foundation models. We show that the existence of these models eliminates the need to train new models for pretext tasks to achieve state-of-the-art performance for SSL-classification. Furthermore, it is comparatively very cheap to use the foundation models instead as not only can we forego training on a pretext task, but we are able to completely bypass the need to train a large convolutional neural network (CNN) or vision transformer (ViT). In fact, we compute the embedding for each of the images in our dataset only one time (as opposed to applying some form of data augmentation during training and making use of the embedding process during training). So, at training time the only weights we need to use at all are those in the model we train on top of the learned representation. This makes our approach incredibly lightweight compared to traditional CNN/ViT training procedures.
4.2 Meta Co-Training
In addition to our method of view construction we propose a novel algorithm for co-training which is based on Meta Pseudo Labels [59], called Meta Co-Training (MCT). Algorithm 1 presents the details of this method.

### v5

@@ Abstract
In many critical computer vision scenarios unlabeled data is plentiful, but labels are scarce and difficult to obtain. As a result, semi-supervised learning which leverages unlabeled data to boost the performance of supervised classifiers have received significant attention in recent literature. One representative class of semi-supervised algorithms are co-training algorithms. Co-training algorithms leverage two different models which have access to different independent and sufficient representations or “views” of the data to jointly make better predictions. Each of these models creates pseudo-labels on unlabeled points which are used to improve the other model. We show that in the common case where independent views are not available, we can construct such views inexpensively using pre-trained models. Co-training on the constructed views yields a performance improvement over any of the individual views we construct and performance comparable with recent approaches in semi-supervised learning. We present Meta Co-Training, a novel semi-supervised learning algorithm, which has two advantages over co-training: (i) learning is more robust when there is large discrepancy between the information content of the different views, and (ii) does not require retraining from scratch on each iteration. Our method achieves new state-of-the-art performance on ImageNet-10% achieving a $\sim 4.7\%$ reduction in error rate over prior work. Our method also outperforms prior semi-supervised work on several other fine-grained image classification datasets.
@@ 1 Introduction
In many critical machine learning scenarios we have access to a large amount of unlabeled data, and relatively few labeled data points for a given task. For standard computer vision (CV) tasks there are large well-known and open source datasets with hundreds of millions or even billions of unlabeled images [47, 46]. In contrast, labeled data is usually an order of magnitude more scarce and otherwise expensive to obtain requiring many human-hours to generate. In this context, semi-supervised learning (SSL) methods are useful as they rely on training more performant models, using small amounts of labeled data and large amounts of unlabeled data.
@@
Not unrelated to, and not to be confused with the idea of semi-supervised learning, is self-supervised learning. [^1] There is also an SSL technique called self-training. Recognizing the unfortunate similarity of these three terms, SSL will always mean semi-supervised learning in this text and we will not abbreviate the other terms. In self-supervised learning a model is trained with an objective that does not require a label that is not evident from the data itself. Self-supervised learning was popularized by the BERT model [18] for generating word embeddings for natural language processing (NLP) tasks. BERT generates its embeddings by solving the pretext task of masked word prediction. Pretext tasks present unsupervised objectives that are not directly related to any particular supervised objective we might want to solve, but rather are solved in the hope of learning a suitable representation to more easily solve a downstream task. These pretext learners are often referred to as foundation models. Several pretext tasks have been proposed for CV to train associated foundation models that solve them [26, 12, 34, 10, 42, 27]. The learned representations for images are often much smaller than the images themselves. As a consequence, this yields the additional benefit of reduced computational cost when using the learned representations.
SSL algorithms [33, 5, 44, 52, 43, 49, 51, 4] involve generating pseudo-labels to serve as weak supervision on unlabeled data and then learning from those pseudo-labels. This weak supervision can either be used to perform consistency regularization, or entropy minimization, or both. Consistency regularization is to enforce that different augmented versions of the same instance are assigned the same label. Entropy minimization is to enforce that all instances are assigned a label with high confidence. Typically consistency regularization is achieved by minimizing a consistency loss between pairs of examples that ought to have the same label [51, 4, 49, 3], and entropy minimization is achieved by iterative re-training [33, 5, 44] or sharpening labels [43, 4, 49, 3].
One particular class of SSL algorithms are co-training [5] algorithms in which two different “views” of the data must be obtained or constructed, and then two different models must be trained and re-trained iteratively. These two views yield models which capture different patterns in their input and thus, informally speaking, are more likely to fail independently. This independence of failure is leveraged so that each model can provide useful labels to the other. Standard benchmark problems in machine learning do not present two views of the problem to be leveraged for co-training. If one hopes to apply co-training, one needs to construct two views from the single view that they have access to. We show that constructing views which satisfy the conditions necessary for co-training is relatively simple. Training different models on the two constructed views can boost performance. Unfortunately, the classical co-training algorithm (CT) fails to utilize pseudo-labels effectively on benchmark tasks.
@@
We propose a novel SSL method which more effectively leverages pseudo-labels. We call our algorithm Meta Co-Training (MCT), because it leverages two views to produce good pseudo-labels as part of a bi-level optimization. In MCT each model is trained to provide better pseudo-labels to the other model, given only its view of the data and the performance of the other model on a labeled set. Similarly to CT, our approach utilizes multiple views to train models which are diverse. These models have an advantage when teaching (and learning from) each-other because they will learn different patterns from their differing input data. We show that our approach provides an improvement over both CT and the current state-of-the-art (SoTA) method [37] on the ImageNet-10% dataset, as well as it establishes new SoTA few-shot performance on several other fine-grained image classification tasks.
Summary of Contributions.
Our contributions are summarized as follows.
@@
    We propose Meta Co-Training, an SSL algorithm which appears to be more robust than co-training and does not require iterative re-training from initialization.
@@
    We establish SoTA top-1 ImageNet-10% [17] accuracy and other few-shot benchmark classification tasks.
@@
    We make our implementation publicly available for others to use and build upon. https://github.com/JayRothenberger/Meta-Co-Training
Outline of the Paper.
In Section 2 we establish notation and provide background information necessary to understand our method. In Section 3 we describe our proposed method, meta co-training. In Section 4 we present our experimental findings on ImageNet [17] and discuss how our method compares to relevant baseline approaches. We perform additional experiments on Flowers102 [41], Food101 [6], FGVCAircraft [39], iNaturalist [29], and iNaturalist 2021 [28] datasets. Additional details such as hyperparameters are provided in Appendix B. In Section 5 we discuss related work with emphasis on semi-supervised methods, including co-training methods. In Section 6 we conclude with a summary and directions for future work.
@@ 2 Background
Notation.
The models that we learn are functions of the form $f\colon\mathcal{X}\rightarrow\Delta^{|\mathcal{Y}|}$ where $\Delta^{|\mathcal{Y}|}$ is the $|\mathcal{Y}|$-dimensional unit simplex. We use $f(x)|_{j}$ to denote the $j$-th value of $f$ in the output. When we want to explicitly refer to a network $f$ parameterized by $\theta$ (e.g., in deep neural networks), we write $f_{\theta}$. For the entirety of this text $\ell$ will refer to the cross-entropy loss which is defined as $\ell\left(y,f_{\theta}(x)\right)=-\sum_{\xi\in\mathcal{Y}}\mathbf{1}\left\{\xi=y\right\}\log\left(f_{\theta}(x)|_{\xi}\right)$. We use $\eta$ as the learning rate and $T$ as the maximum number of steps (or updates) in a gradient-based optimization procedure. Finally, we use $\mathbf{1}\left\{\mathcal{A}\right\}$ as an indicator function of an event $\mathcal{A}$; that is, $\mathbf{1}\left\{\mathcal{A}\right\}$ is equal to 1 when $\mathcal{A}$ holds, otherwise 0.
Semi-Supervised Learning.
@@
We are interested in situations where we have a large pool of unlabeled instances $U$ as well as a small portion $L$ of them that is labeled. That is, the learning algorithm has access to a dataset $S=L\cup U$, such that $L=\left\{(x_{i},y_{i})\right\}_{i=1}^{m}$ and $U=\left\{x_{j}\right\}_{j=1}^{u}$. For $L$ the instance part is denoted as $X_{L}$ and the label part is denoted as $Y_{L}$. In this setting, semi-supervised learning (SSL) methods attempt to first learn an initial model $f_{init}$ using the labeled data $L$ via a supervised learning algorithm, and in sequence, an attempt is being made in order to harness the information that is hidden in the unlabeled set $U$, so that a better model $f$ can be learnt.
View Construction.
A view, in the context of co-training methods, is a subset of the input features. A view typically has two properties: (i) it is sufficient for predicting the desired quantity, and (ii) it is conditionally independent of other views given the label. Previous methods of view construction for co-training include manual feature subsetting [19], automatic feature subsetting [11], random feature subsetting [7], random subspace selection [8], and adversarial examples [44]. We use self-supervised learning to construct views for our experiments. However, if one has access to a dataset where the instances have two different views, then one can still obtain useful representations with this approach using the same self-supervised model or none at all.
While our approach is more widely applicable, it is particularly well-suited to computer vision. There are a plethora [23, 12, 10, 34, 45, 42, 26] of competitive learned representations for images. Choosing foundation model representations as views makes our approach lightweight enough to be feasible with limited hardware resources.
3 Proposed Method: Meta Co-Training
We propose a novel semi-supervised learning algorithm called Meta Co-Training (MCT), which is described below. MCT operates within a bi-level student-teacher optimization framework. A student is optimized to replicate the labels its teacher gives to a set of unlabeled points while a teacher is optimized to provide labels to the same unlabeled points such that its student learns to predict well on labeled data. We call MCT a co-training algorithm because it operates on two views of the data, however unlike the original co-training method [5] we do not train multiple models for each view sequentially. Each view has exactly one model which acts as both a teacher for the other view and a student for its own view. That is to say that the model learning on view 1 will take an unlabeled instance from view 1 as input and provide the pseudo-label prediction to the student for the same unlabeled instance in view 2. The student – the model learning on view 2 – will use the pseudo-label to learn to predict similarly.
The lower of the two levels of the optimization is the student optimization (Equation 2). The student parameters $\theta_{S}$ are optimized as a function of the teacher parameters $\theta_{T}$:
$\mathcal{L}_{u}\left(\theta_{T},\theta_{S}\right)=\ell\left(\operatorname*{arg\,max}_{\xi}f_{\theta_{T}}(U)|_{\xi},f_{\theta_{S}}(U)\right)$ (1)
$\theta_{S}^{\prime}=\theta_{S}^{\text{PL}}(\theta_{T})\in\operatorname*{arg\,min}_{\theta_{S}}\mathcal{L}_{u}\left(\theta_{T},\theta_{S}\right)\,.$ (2)
The teacher optimization (Equation 4) forms the upper level of the optimization:
$\mathcal{L}_{L}\left(\theta_{S}^{\prime}\right)=\ell\left(Y_{L},f_{\theta_{S}^{\prime}}(X_{L})\right)$ (3)
$\theta_{T}^{\prime}=\operatorname*{arg\,min}_{\theta_{T}}\displaystyle\mathcal{L}_{L}\left(\theta_{S}^{\prime}\right)\,.$ (4)
All together the objective of MCT from the perspective of the current view model as the teacher is
$\min_{\theta_{T}}\mathcal{L}_{u}\left(\theta_{T},\theta_{S}\right)+\mathcal{L}_{L}\left(\theta_{S}^{\prime}\right)\,.$ (5)
It is cumbersome to speak of teachers and students when each model is both. We give the objective of MCT in Equation 6 as jointly optimizing the following objectives with $f_{\theta_{1}}$ the model learning on view 1 and $f_{\theta_{2}}$ the model learning on view 2.
$\min_{\theta_{1},\theta_{2}}\mathcal{L}_{u}\left(\theta_{1},\theta_{2}\right)+\mathcal{L}_{L}\left(\theta_{2}^{\prime}\right)\ +\mathcal{L}_{u}\left(\theta_{2},\theta_{1}\right)+\mathcal{L}_{L}\left(\theta_{1}^{\prime}\right)\,$ (6)
Co-training algorithms are traditionally evaluated using the joint prediction of the constituent models; i.e., the re-normalized element-wise product of the individual predictions.
Figure 1: At each step $t\in\{1,\ldots,T\}$ of meta co-training the models that correspond to the so-far learnt parameters $\theta_{1;t}$ and $\theta_{2;t}$ play the role of the student and the teacher simultaneously using batches for their respective views. Pseudo-labeling occurs on complementary views so that the teacher can provide the student with labels on an unlabeled batch. Labeled batches may, or may not, use complementary views as the purpose that they serve is to calculate the risk of the student model on the labeled batch and this result signals the teacher model to update its weights accordingly.
To make tractable the optimization in Equation 6 we approximate the student and teacher optimizations using alternating iterations of stochastic gradient descent. In Algorithm 1 the student updates are in lines 13-14 while the teacher updates are in lines 24-25. An overview of MCT is provided in Figure 1. At each step $t\in\{1,\ldots,T\}$ of meta co-training the models that correspond to the so-far learnt parameters $\theta_{1;t}$ and $\theta_{2;t}$ participate in a student-teacher framework in which each model plays the role of the student for their view and the teacher for the other model simultaneously. The representations obtained from the view construction process of Section 2 form different views that are used by MCT.
By training two models on two views to collaborate we can take advantage of the aspects of MPL and Co-Training that make each algorithm successful. Co-training is more effective when the two constituent models fail (predict incorrectly) independently. During MCT the teacher model will receive positive feedback precisely when it makes a prediction that the other model would not have made which results in an improvement. MCT trains models in a way which encourages each teacher to express independence while it teaches each student to incorporate the improvement. Naturally, once the two models predict similarly then the gradient of the student parameters with respect to the teacher’s labels is zero and they both cease to improve. This, however, is also when MPL ceases to improve and it is unlikely to have models which are as diverse as those which are trained from independent views.
Algorithm 1 Meta Co-Training (Section 2 has notation)